OpenStudy (anonymous):

multiply 3√3 * 3√9 = 3 answer pls check for me

OpenStudy (anonymous):

No that's not right, √9 = 3, so we have 3√3*3*3=27√3

OpenStudy (anonymous):

Thomas 9 could you show how you arrive at this ans. pls

OpenStudy (anonymous):

\[\sqrt{9}=3\] right? So: \[3\sqrt{3}\cdot 3\sqrt{9}=3\sqrt{3}\cdot3\cdot 3=3\sqrt{3}\cdot9=9\cdot 3\sqrt{3}=27\sqrt{3}\]

OpenStudy (anonymous):

ok thnks